OpenStudy (anonymous):

simplify the square root of 27x to the fifth power

OpenStudy (anonymous):



OpenStudy (anonymous):

IS that the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

 So, 27=3*3*3=3*3^2